# When Hyperspectral Image Classification Meets Diffusion Models: An Unsupervised Feature Learning Framework

### v1

@@ IV-C Quantitative Results and Analysis
((e))
((f))
((g))
((h))
((i))
((j))
Figure 4: Classification maps obtained by different methods on the Indian Pines dataset. (a) Ground truth. (b) 2-D CNN (OA=87.77%). (c) 3-D CNN (OA=85.42%). (d) SSRN (OA=97.75%). (e) SF (OA=92.31%). (f) SSFTT (OA=97.47%). (g) GAHT (OA=97.95%). (h) 3DCAE (OA=92.69%). (i) UMSDFL (OA=97.02%). (j) Diff-HSI (OA=99.46%).
@@
((e))
((f))
((g))
((h))
((i))
((j))
Figure 5: Classification maps obtained by different methods on the Pavia University dataset. (a) Ground truth. (b) 2-D CNN (OA=98.86%). (c) 3-D CNN (OA=97.88%). (d) SSRN (OA=99.10%). (e) SF (OA=97.54%). (f) SSFTT (OA=99.21%). (g) GAHT (OA=99.53%). (h) 3DCAE (OA=96.77%). (i) UMSDFL (OA=99.02%). (j) Diff-HSI (OA=99.94%).
@@
((d))
((e))
((f))
((g))
((h))
((i))
((j))
Figure 6: Classification maps obtained by different methods on the Houston2018 dataset. (a) Ground truth. (b) 2-D CNN (OA=94.84%). (c) 3-D CNN (OA=86.88%). (d) SSRN (OA=91.61%). (e) SF (OA=90.65%). (f) SSFTT (OA=95.48%). (g) GAHT (OA=96.69%). (h) 3DCAE (OA=90.34%). (i) UMSDFL (OA=95.38%). (j) Diff-HSI (OA=98.07%).
IV-D Ablation Studies
In this section, we analyze the effect of the components in our method.
IV-D1 Ablation for Set Number of Dynamic Weights
Our dynamic feature fusion module contains $n$ sets of dynamic weights which impact the classification performance. The change of results by changing $n$ is shown in Tabel VII. We reveal that the best performance is achieved when $n$ is taken as 3. With smaller $n$, only part of the useful information is obtained from the feature bank, which is insufficient for classification. On the other hand, larger $n$ will bring more parameters and excess information.
IV-D2 Ablation for Pretraining Steps
We validate the effectiveness of pretraining on the Indian Pines dataset in terms of OA, AA, and $\kappa$. As shown in Table VIII, only 10k steps pretraining brings dramatic improvement (more than 30% OA) to the final classification performance. Furthermore, as the pretraining steps increase, the performance continues to rise to the best at around 40k steps.
IV-D3 Sensitivity Analysis of Timestep T
To analyze the features extracted from different timestep $t$, we record the change of the classification performance when changing $t$. For easy understanding, we choose 4 of the 16 classes to show their changes, as illustrated in Fig. 7. The performance for each class behaves differently as the $t$ increases. Features of larger $t$ are more sensitive to class ”corn-notill” and features of smaller $t$ are more informative to class ”woods”. For class ”Grass-pasture-mowed” and class ”corn”, features extracted at the intermediate $t$ is the most discriminative. Thus, an appropriate set of $t$ is vital for accurate performance.
TABLE VII: Ablation analysis of the proposed dynamic feature fusion module with a different number of dynamic weight sets in terms of OA, AA, and $\kappa$ on the Indian Pines dataset.
Metric The number of the sets of dynamic weights $n$
1 2 3 4
OA (%) 99.36 99.38 99.46 99.43
AA (%) 99.20 99.20 99.37 99.28
$\kappa$ 0.9918 0.9930 0.9938 0.9935
TABLE VIII: Effect of Pretraining Steps in terms of OA, AA, and $\kappa$ on the Indian Pines dataset.
Metric The number of pretraining steps
0 10k 20k 30k 40k 50k
OA (%) 68.04 97.53 98.40 99.39 99.46 99.45
AA (%) 48.95 96.64 97.76 99.30 99.37 99.39
$\kappa$ 58.51 0.9718 0.9817 0.9930 0.9938 0.9936
TABLE IX: Classification performance with different sets of timestep $t$ on the Indian Pines dataset.
Timestep Sets $X$ Metric
OA (%) AA (%) $\kappa$
$[100,200,300]$ 98.81 98.72 0.9874
$[700,800,900]$ 98.77 98.68 0.9872
$[100,500,900]$ 98.87 98.83 0.9886
Dynamic Feature Fusion 99.46 99.37 0.9938
Table IX shows the results with several sets of manually assigned $t$. Specifically, given a timestep set $X=[t_{1},t_{2},t_{3}]$, we extract the features at timestep $t_{1}$, $t_{2}$, and $t_{3}$ to construct the pixel-wise representation for classification. As Table IX presents, the performance is sensitive to the choice of $t$. $X$ with a large $t$, a small $t$ and a medial $t$ performs better than $X$ with only smaller $t$ or larger $t$. Because the former contains a wider range of $t$ leading to more diverse features which make the representation more general for samples of all classes. However, representation obtained by the manually designed method is not always the most appropriate for each sample, since it sacrifices some specificity to improve generalization. So we propose the dynamic feature fusion module to customize $T$ for each sample which makes it more accurate for classification. The results in Table IX demonstrate that our dynamic feature fusion module learns more effective representation compared with the manually selected method.
@@ IV-D3 Sensitivity Analysis of Timestep T
Figure 8: Classification results of different patch sizes on the Indian Pines, Pavia University, and Houston2018 Dataset.
Figure 9: Classification results of different PCA components on the Indian Pines, Pavia University, and Houston2018 Dataset.
IV-E Parameter Analysis
@@
In this section, we analyze the effect of various parameters that influence classification performance by training our proposed Diff-HSI in the same experimental setting as Section IV-B with different parameters.
IV-E1 Effect of Different Patch Size
@@
First, we discuss the effect of different patch sizes on classification performance by two indexes, OA and AA. As shown in Fig. 8, the patch size varies from 24$\times$24 to 60$\times$60. as the patch size increases, the performance first increases and then decreases. The best performance is obtained when the patch size is 48$\times$48 with OA of 99.39% and AA of 99.30%. Too small patches contain insufficient spatial information and too large patches reduce the attention to detailed structures. Thus, we choose 48$\times$48 to be the patch size for the proposed Diff-HSI.
IV-E2 Effect of Different PCA Components
@@
This section analyzes the influence of the number of PCA components, which determines how much spectral information is retained in the compressed data. As the number of PCA components increases, more spectral information is retained while more computational cost and more redundant information are brought. Since each dataset has a different number of channels, the range of PCA components is different for the three datasets. Assuming that $N$ is the channel number of a dataset, the number of PCA components varies from $N/20$ to $N/5$. According to the results shown in Fig. 9, the best performance is achieved at PCA components of $N/8$.
@@ V Conclusion
HSI contains rich spectral-spatial information and complex relations, which are critical for classification tasks. We propose the Diff-HSI method to learn discriminative spectral-spatial features in an unsupervised manner using the diffusion model. Most of the existing works perform HSI classification in an explicit way based on CNN or Transformer. These methods often heavily rely on the number of labeled data and are insufficient to model complex relations, limiting the improvement in performance. Some works explore unsupervised learning for HSI classification through an encoder-decoder feature learning method. However, limited by the feature-extracting capability of the model, they mainly tend to capture low-level features, lacking high-level features which are significant for discriminative spectral-spatial feature representation. Our proposed method leverages the powerful implicit learning ability of the diffusion model to learn effective spectral-spatial features with high degrees of freedom, constructing the step-wise feature bank. Furthermore, we design a dynamic feature fusion model to obtain multi-timestep representation from the feature bank automatically. Quantitative experiments on three HSI datasets demonstrate that our proposed Diff-HSI outperforms state-of-art supervised and unsupervised methods.

### v2

@@ IV-D1 Ablation for Class & Timestep-oriented Multi-Stage Feature Purification
Figure 9: Classification performance of the compared methods with different proportions of training samples on four datasets. (a) Indian Pines. (b) PaviaU. (c) Houston 2018. (d) Longkou.
TABLE IX: Ablation for Class & Timestep-oriented Multi-Stage Feature Purification (CTMSFP) on four datasets in terms of OA, inference time (IT), parameters and GPU memory. The best result are shown in bold.
Method Indian Pines PaviaU Houston 2018 Longkou Param. (M) GPU memory (G)
OA (%) IT (s) OA (%) IT (s) OA (%) IT (s) OA (%) IT (s)
SpectralDiff 98.54 10.92 99.46 22.41 95.28 268.88 98.71 178.25 1.38 3.09
MTMSD (w/o CTMSFP)
99.39 4.55 99.90 22.78 98.20 215.20 99.51 85.14 55.91 1.28
MTMSD 99.45 3.52 99.95 12.41 98.29 168.52 99.61 61.53 20.18 0.46
IV-D2 Ablation for Selective Timestep Feature Fusion
The ablation results for the selective timestep feature fusion are shown in Table. X. In the manual selection method, we extract features from a single timestep for classification. The optimal timestep is selected for each dataset, and the best result is shown in the table. However, using the features from only one single timestep leads to the loss of abundant spectral-spatial information. Although the optimal timestep is chosen for each dataset, it lacks adequate information, only containing textural features or semantics to model spectral-spatial relations, and is not flexible enough to accommodate different patch data, both of which limit performance. The average fusion method considers features from different timestep $t$, obtaining better results than manual selection. However, assigning a uniform selection weight to features across all the timesteps indiscriminately does not fully harness their potential because the significance of features from different $t$ for different data instances is heterogeneous. Therefore, the proposed selective fusion is more optimal for feature fusion. Compared with the average fusion, our selective timestep feature fusion achieves better performance by assigning the vital timestep features higher selecting weights. Furthermore, compared with no global-feature guidance, our proposed selective timestep fusion with global-feature guidance improves further due to the supplement of global information that enhances the ability to represent spatial distributions.
TABLE X: OA (%) of the proposed MTMSD with different feature fusion and global-feature guidance on four datasets. the best results are shown in bold.
Method Guidance Indian Pines PaviaU Houston 2018 Longkou
Manual Selection 98.17 99.40 94.57 99.08
Average Fusion 99.03 99.78 97.85 99.39
Selective Fusion 99.31 99.90 98.07 99.53
99.45 99.95 98.29 99.61
IV-E Discussion and Visualization
IV-E1 Analysis of Features from Multiple Timesteps
To analyze the features extracted from different timestep $t$, we record the change of the classification performance when changing $t$. For easy understanding, we choose 4 of the 16 classes to show their changes, as illustrated in Fig. 10. The performance for each class behaves differently as the $t$ increases. Features of larger $t$ are more sensitive to class ”corn-notill” and features of smaller $t$ are more informative to class ”woods”. For class ”Grass-pasture-mowed” and class ”corn”, features extracted at the intermediate $t$ is the most discriminative. Thus, an appropriate fusion of features at different $t$ is vital for accurate performance.
((a))
((b))
Figure 11: Feature visualization of retained and removed channels after class & timestep-oriented multi-stage purification at different timestep $t$. (a) Retained. (b) Removed.
@@ IV-E1 Analysis of Features from Multiple Timesteps
Figure 12: Feature visualization of different timestep $t$ with higher and lower selection weights. (a) Pseudocolor images. (b) Ground truth. (c) Higher weights. (d) Lower weights.
@@
Figure 13: Classification results of different parameters on the Indian Pines, PaviaU, Houston 2018, and Longkou Dataset. (a) Patch size. (b) PCA component. (c) Pretraining Steps.
IV-E2 Visualization of Multi-Timestep Multi-Stage Diffusion Feature Exploration
The feature map of retained and removed channels after class & timestep-oriented multi-stage purification are visualized in Fig. 11. The retained channels capture more semantic information compared to the removed ones after purification, as evident from Fig. 11. Additionally, for the same channel, features at different timesteps exhibit diversity. Features at shallow timesteps tend to capture stochastic details, while those at larger timesteps focus more on higher-level semantic information.
To further demonstrate the selective timestep feature fusion’s ability to select features at different timesteps, we visualized the features with higher and lower selection weights for the same sample. As shown in Fig. 12, the results indicate that for features with higher selection weights, the response of the target pixels is more consistent with surrounding pixels of the same class, containing more classification-correlated information. If the features across all timesteps are averaged without discrimination, the important information relevant to the classification task may become obfuscated. Thus, to obtain an effective representation suitable for hyperspectral image classification, our proposed selective timestep feature fusion increases the proportion of classification-related information in the representation by assigning higher weights to more discriminative features. The visualization provides evidence that our proposed selective timestep feature fusion can effectively select timestep features suitable for classification tasks.
IV-F Parameter Analysis
@@
In this section, we analyze the effect of various parameters that influence classification performance by training our proposed MTMSD in the same experimental setting as Section IV-B with different parameters.
IV-F1 Effect of Different Patch Size
@@
First, we discuss the effect of different patch sizes on classification performance by two indexes, OA and AA. As shown in Fig. 13(a), the patch size varies from 24$\times$24 to 60$\times$60. As the patch size increases, the performance first increases and then decreases. The best performance is obtained when the patch size is 48$\times$48 with OA of 99.39% and AA of 99.30%. Too small patches contain insufficient spatial information and too large patches reduce the attention to detailed structures. Thus, we choose 48$\times$48 to be the patch size for the proposed MTMSD.
IV-F2 Effect of Different PCA Components
@@
This section analyzes the influence of the number of PCA components, which determines how much spectral information is retained in the compressed data. As the number of PCA components increases, more spectral information is retained while more computational cost and more redundant information are brought. Since each dataset has a different number of channels, the range of PCA components is different for four datasets. Assuming that $N$ is the channel number of a dataset, the number of PCA components varies from $N/20$ to $N/5$. According to the results shown in Fig. 13(b), the best performance is achieved at PCA components of $N/8$.
IV-F3 Effect of Different Pretraining Steps
We validate the effectiveness of pretraining on the four datasets in terms of OA. As shown in Fig. 13(c), only 10k steps pretraining brings dramatic improvement (more than 30% OA) to the final classification performance. Furthermore, as the pretraining steps increase, the performance continues to rise to the best at around 40k steps.
IV-G Efficiency Analysis
We evaluate the inference time of different methods on four public datasets to analyze the efficiency. All the experiments are carried out on a Nvidia RTX 3090. As shown in Table XI, although our method is not the fastest in inference speed, it achieves the best classification performance since it utilizes effective multi-timestep multi-stage diffusion features. It is noted that our method has reduced the average inference time by 62% across four datasets compared to another diffusion-based HSI classification method, SpectralDiff.
TABLE XI: GPU inference time(s) of different methods.
Method Indian Pines PaviaU Houston 2018 Longkou
2-D CNN 3.61 7.02 47.47 69.86
3-D CNN 3.83 7.66 48.99 65.24
SSRN 2.74 5.36 35.94 37.99
SF 3.52 7.59 63.73 48.40
SSFTT 3.79 7.28 31.38 59.32
GAHT 2.51 5.25 44.90 34.19
3DCAE 6.12 12.06 193.07 83.91
3DAES 2.81 5.64 41.64 41.83
UMSDFL 14.28 31.04 293.41 249.03
SpectralDiff 10.92 22.41 268.88 178.25
MTMSD (Ours) 3.52 12.41 168.52 61.53
@@ V Conclusion
HSI contains rich spectral-spatial information and complex relations, which are critical for classification tasks. Many supervised and unsupervised deep learning methods are proposed to learn spectral-spatial features from HSI data, achieving promising results in HSI classification. Recently, diffusion models as powerful models in generation and reconstruction tasks have been applied to HSI classification in one recent work. However, the diffusion features used in the work are extracted solely from a single timestep and a single stage of the denoising U-Net manually selected for each dataset, which limits the performance. Thus, we propose a diffusion-based feature learning framework that explores Multi-Timestep Multi-Stage Diffusion features for HSI classification for the first time, named MTMSD. Quantitative experiments on four HSI datasets demonstrate that our proposed MTMSD outperforms state-of-the-art supervised and unsupervised methods.